Q is the midpoint of segment PR. What are PQ, QR and PR?

PQ = 6x - 7

QR = 5x + 1


1
Expert's answer
2021-09-13T17:13:06-0400

PQ = QR,

6x - 7 = 5x + 1,

x = 8,

PQ = QR = 5×8 + 1 = 41,

PR = 41×2 = 82.